Solve the porportion


14.d/6 = 65/26
Any help with this equation would be Greatly appreciated.

Answer by checkley71(8403)

D/6=65/26 CROSSS MULTIPLY
26D=6*65
26D=390
D=390/26
D=15 ANSWER.
PROOF:
15/6=65/26
15*26=6*65
390=390